# Bayesian Density-Density Regression with Application to Cell-Cell Communications

### v1

@@ 6.2 Semi-Simulation
We validate the proposed graph discovery framework by setting up a simulation using simulation truths inspired by the OneK1K data. Particularly, we use the real predictor distributions taken from the data and a simulation truth for $f_{A,b}$ estimated from the real data to set up a simulation truth for the response distributions. Assume then from the earlier data analysis we have a posterior Monte Carlo sample with $T$ posterior samples under the Bayesian DDR model for all pairs of cell types. Recall that $\widetilde{d}_{e}\in\{0,1\}$ is an indicator for including edge $e$ in the inferred graph under Bayesian DDR (Figure 4c). We define the following empirical distributions:
@@
Under scenario 1 (“no edge”) we sample for each pair $e=(\ell,r)$ of cell types the simulation truth $A_{eij}^{\star}\sim p_{0}$ and $b_{ej}^{\star}\sim p_{0}$ ($i=1,\ldots,d_{e1},~{}j=1,\ldots,d_{e2}$). Under scenario 2 (“full graph”) we sample the simulation truth for all pairs $e=(\ell,r)$ as $A_{eij}^{\star}\sim p_{1}$ and $b_{ej}^{\star}\sim p_{1}$ ($i=1,\ldots,d_{e1}),~{}j=1,\ldots,d_{e2}$). Finally, under scenario 3 (“sparse graph”) we use the inferred graph under Bayesian DDR (Figure 4b) as the ground truth. For included edges $e$, we sample the simulation truth like under scenario 2. For excluded edges we sample like under scenario 1. Under all three scenarios, after generating the simulation truths $(A_{e},b_{e})$ for the linear mappings, we simulate response distributions ${G}_{ei}=\frac{1}{m_{ei}}\sum_{j=1}^{m_{ei}}\delta_{A^{\star}_{e}x_{ij}+b_{e}^{\star}+\epsilon_{0}}$ where $x_{ij}$ is the observed ligands gene expression from the $i$-th donnor in the OneK1K data, and $\epsilon_{0}\sim\mathcal{N}(0,0.01I)$ for $i=1,\ldots,N_{e}$. This concludes the definition of the simulation truth. We then implement inference under the Bayesian DDR and the Bayesian MLR model, with the same hyperparameters as in the earlier data analysis. We then use FDR and FNR control to report a graph.
Inference under Bayesian DDR. Figure 5 shows the simulated true graphs and the reported graphs under the Bayesian DDR model for all three scenarios. Under scenario 1 (“no edge”) we find the optimal threshold as $\epsilon=1$, yielding $(\overline{\text{FDR}},\overline{\text{FNR}})=(0.035,0.064)$, and 5 edges are found. Under scenario 2 (“full graph”), $\epsilon=1$ with $(\overline{\text{FDR}},\overline{\text{FNR}})=(0.083,0.125)$, and 7 edges are discovered. Finally, under scenario 3 (“sparse graph”), $\epsilon=1$, and $(\overline{\text{FDR}},\overline{\text{FNR}})=(0.036,0.17)$. In this setup, Bayesian DDR yields 9 edges, with 6 correct edges and 3 incorrect edges. Boxplots of RPE for all ordered pairs of cell types, and under all three scenarios is shown in in Figure 23 in Supplementary Materials C. The relatively low RPE values show a good fit for almost all ordered pairs of cell types across all three settings. Additionally, Figure 24 in Supplementary Materials C shows boxplots of squared errors for the parameters ($A_{e},b_{e}$) for the three settings and all pairs of cell types. Overall we find relatively small parameter errors in all cases.
@@
The previous simulation included a linear (in the atoms) simulation truth, favoring MLR (since it also implies linearity of mean bulk gene expression). To highlight the limitations of linear regression for bulk gene expression $(\bar{x},\bar{y})$ we consider another simulation, with: ${F}_{i}=\frac{1}{m}\sum_{j=1}^{m}\delta_{x_{ij}},\quad{G}_{i}=\frac{1}{m}\sum_{j=1}^{m}\delta_{(A^{\star}x_{ij}+b^{\star})^{2}+\epsilon_{0}},$ where $A^{\star}=\left[\begin{smallmatrix}1&0\\ 1&1\end{smallmatrix}\right]$, $b^{\star}=[1,1]^{\prime}$, $\epsilon_{0}\sim\mathcal{N}([0,0]^{\prime},0.01I)$, and $m=100$. We sample $x_{i1},\ldots,x_{im}\overset{i.i.d.}{\sim}\mathcal{N}(\mu_{i},\Sigma_{i})$ with $\mu_{i}\sim\mathcal{N}([0,0]^{\prime},I)$ and $\Sigma_{i}\sim IW(10,I)$ for $i=1,\ldots,N=10$. We implement Bayesian DDR with $f_{A,b}(x)=(Ax+b)^{2}$ (element-wise quadratic), and compare with inference under a Bayesian multiple nonlinear regression (MNR) assuming $\bar{y}\mid\bar{x},A,\Sigma\sim\mathcal{MN}((\bar{x}A)^{2},\Sigma,I)$ (element-wise quadratic). Figure 6 shows the fitted densities under Bayesian DDR and fitted means under the Bayesian MNR model. While inference under the Bayesian DDR can still capture the high-density regions, Bayesian MNR naturally gives inaccurate fits of the mean, due to model mis-specification. Similarly, in cases where $\mathbb{E}[X_{i}]=0$ or $\mathbb{E}[X_{i}]$ does not exist (e.g., Cauchy distribution), Bayesian MLR based on the mean statistic is not feasible, while Bayesian DDR remains well-defined.
@@ 7 Conclusion
We introduced Bayesian Density-Density Regression (Bayesian DDR) as a novel framework for modeling complex regression relationships between multivariate random distributions.
@@
Future work will focus on addressing these limitations. For instance, we may extend the framework to Bayesian nonparametric density-density regression. In contrast to recent frequentist nonparametric density-density regression approaches (Chen et al., , 2023; Ghodrati and Panaretos, , 2022), which focus on one-dimensional regression, we envision a Bayesian nonparametric prior for the regression function in multivariate settings. Finally, one could explore alternative geometric variants of the SW distance for the generalized likelihood, and other loss functions or policies for graph discovery.
@@
Supplementary Materials for “Bayesian Density-Density Regression with Application to Cell-Cell Communications”
Appendix A MCMC

### v2

@@ 6.2 Semi-Simulation
We validate the proposed graph discovery framework by setting up a simulation using simulation truths inspired by the OneK1K data. Particularly, we use the real predictor distributions taken from the data and a simulation truth for $f_{A,b}$ estimated from the real data to set up a simulation truth for the response distributions. Assume then that we have a posterior Monte Carlo sample with $T$ posterior samples under the Bayesian DDR model for all pairs of cell types. Recall that $\widetilde{d}_{e}\in\{0,1\}$ is an indicator for including edge $e$ in the inferred graph under Bayesian DDR (Figure 5c). We define the following empirical distributions:
@@
Under scenario 1 (“no edge”) we sample for each pair $e=(\ell,r)$ of cell types the simulation truth $A_{eij}^{\star}\sim p_{0}$ and $b_{ej}^{\star}\sim p_{0}$ ($i=1,\ldots,d_{e1},~j=1,\ldots,d_{e2}$). Under scenario 2 (“full graph”) we sample the simulation truth for all pairs $e=(\ell,r)$ as $A_{eij}^{\star}\sim p_{1}$ and $b_{ej}^{\star}\sim p_{1}$ ($i=1,\ldots,d_{e1}),~j=1,\ldots,d_{e2}$). Finally, under scenario 3 (“sparse graph”) we use the inferred graph under Bayesian DDR (Figure 5b) as the ground truth. For included edges $e$, we sample the simulation truth like under scenario 2. For excluded edges we sample like under scenario 1. Under all three scenarios, after generating the simulation truths $(A_{e},b_{e})$ for the linear mappings, we simulate response distributions ${G}_{ei}=\frac{1}{m_{ei}}\sum_{j=1}^{m_{ei}}\delta_{A^{\star}_{e}x_{ij}+b_{e}^{\star}+\epsilon_{0}}$ where $x_{ij}$ is the observed ligands gene expression from the $i$-th donnor in the OneK1K data, and $\epsilon_{0}\sim\mathcal{N}(0,0.01I)$ for $i=1,\ldots,N_{e}$. This concludes the definition of the simulation truth. We then implement inference under the Bayesian DDR and the Bayesian MLR model, with the same hyperparameters as in the earlier data analysis. Finally, we use FDR and FNR control to report a graph.
Inference under Bayesian DDR. Figure 6 shows the simulated true graphs and the reported graphs under the Bayesian DDR model for all three scenarios. Under scenario 1 (“no edge”) we find the optimal threshold as $\epsilon=1$, yielding $(\overline{\text{FDR}},\overline{\text{FNR}})=(0.035,0.064)$, and 5 edges are reported. Under scenario 2 (“full graph”), $\epsilon=1$ with $(\overline{\text{FDR}},\overline{\text{FNR}})=(0.083,0.125)$, and 7 edges are discovered. Finally, under scenario 3 (“sparse graph”), $\epsilon=1$, and $(\overline{\text{FDR}},\overline{\text{FNR}})=(0.036,0.17)$. In this setup, Bayesian DDR yields 9 edges, with 6 correct edges and 3 incorrect edges. Boxplots of RPE for all ordered pairs of cell types, and under all three scenarios is shown in in Figure 24 in Supplementary Materials D. The relatively low RPE values show a good fit for almost all ordered pairs of cell types across all three settings. Additionally, Figure 25 in Supplementary Materials D shows boxplots of squared errors for the parameters ($A_{e},b_{e}$) for the three settings and all pairs of cell types. Overall we find relatively small parameter errors in all cases.
@@
The previous simulation included a linear (in the atoms) simulation truth, favoring MLR (since it also implies linearity of mean bulk gene expression). To highlight the limitations of linear regression for bulk gene expression $(\bar{x},\bar{y})$ we consider another simulation, with: ${F}_{i}=\frac{1}{m}\sum_{j=1}^{m}\delta_{x_{ij}},\quad{G}_{i}=\frac{1}{m}\sum_{j=1}^{m}\delta_{(A^{\star}x_{ij}+b^{\star})^{2}+\epsilon_{0}},$ where $A^{\star}=\left[\begin{smallmatrix}1&0\\ 1&1\end{smallmatrix}\right]$, $b^{\star}=[1,1]^{\prime}$, $\epsilon_{0}\sim\mathcal{N}([0,0]^{\prime},0.01I)$, and $m=100$. We sample $x_{i1},\ldots,x_{im}\overset{i.i.d.}{\sim}\mathcal{N}(\mu_{i},\Sigma_{i})$ with $\mu_{i}\sim\mathcal{N}([0,0]^{\prime},I)$ and $\Sigma_{i}\sim IW(10,I)$ for $i=1,\ldots,N=10$. We implement Bayesian DDR with $f_{A,b}(x)=(Ax+b)^{2}$ (element-wise quadratic), and compare with inference under a Bayesian multiple nonlinear regression (MNR) assuming $\bar{y}\mid\bar{x},A,\Sigma\sim\mathcal{MN}((\bar{x}A)^{2},\Sigma,I)$ (element-wise quadratic). Figure 7 shows the fitted densities under Bayesian DDR and fitted means under the Bayesian MNR model. While inference under the Bayesian DDR can still often capture the high-density regions, Bayesian MNR naturally gives inaccurate fits of the mean, due to model mis-specification. Similarly, in cases where $\mathbb{E}[X_{i}]=0$ or $\mathbb{E}[X_{i}]$ does not exist (e.g., Cauchy distribution), Bayesian MLR based on the mean statistic is not feasible, while Bayesian DDR remains well-defined.
@@ 7 Conclusion
e introduced Bayesian Density-Density Regression (Bayesian DDR) as a novel framework for modeling complex regression relationships between multivariate random distributions.
@@
Future work will focus on addressing these limitations. For instance, we may extend the framework to Bayesian nonparametric density-density regression. In contrast to recent frequentist nonparametric density-density regression approaches (Chen et al.,, 2023; Ghodrati and Panaretos,, 2022), which focus on one-dimensional regression, we envision a Bayesian nonparametric prior for the regression function in multivariate settings. One could explore alternative geometric variants of the SW distance for the generalized likelihood, and other loss functions or policies for graph discovery. Finally, to deal with a more diverse set of distribution, Orlic-Wasserstein (Sturm,, 2011) can be considered for the generalized likelihood.
@@
Supplementary Materials for “Bayesian Multivariate Density-Density Regression”
Appendix A Proofs
A.1 Proof of Theorem 1
Lemma 1.
For any $G_{1}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $G_{2}\in\mathcal{P}_{2}(\mathbb{R}^{d})$, we have:
$SW_{2}(G_{1},G_{2})\leq W_{2}(G_{1},G_{2}).$ (23)
Proof.
From the definitions, we have:
$SW_{2}^{2}(G_{1},G_{2})$ $=\mathbb{E}_{\theta\sim\mathcal{U}(\mathbb{S}^{d-1})}[W_{2}^{2}(\theta\sharp G_{1},\theta\sharp G_{2})]$ (24)
$=\mathbb{E}_{\theta\sim\mathcal{U}(\mathbb{S}^{d-1})}\left[\inf_{\pi\in\Pi(G_{1},G_{2})}\mathbb{E}_{(X,Y)\sim\pi}[(\theta^{\top}X-\theta^{\top}Y)^{2}]\right]$ (25)
$=\mathbb{E}_{\theta\sim\mathcal{U}(\mathbb{S}^{d-1})}\left[\inf_{\pi\in\Pi(G_{1},G_{2})}\mathbb{E}_{(X,Y)\sim\pi}[(\theta^{\top}(X-Y))^{2}]\right]$ (26)
$\leq\mathbb{E}_{\theta\sim\mathcal{U}(\mathbb{S}^{d-1})}\left[\inf_{\pi\in\Pi(G_{1},G_{2})}\mathbb{E}_{(X,Y)\sim\pi}[\|\theta\|_{2}^{2}\|X-Y\|_{2}^{2}\right]\quad\text{(Cauchy–Schwarz)}$ (27)
$=\mathbb{E}_{\theta\sim\mathcal{U}(\mathbb{S}^{d-1})}\left[\inf_{\pi\in\Pi(G_{1},G_{2})}\mathbb{E}_{(X,Y)\sim\pi}[\|X-Y\|_{2}^{2}\right]\quad(\|\theta\|_{2}^{2}=1)$ (28)
$=W_{2}^{2}(G_{1},G_{2}),$ (29)
which completes the proof. ∎
Lemma 2.
Under Assumptions 3 and 5, we have
$\big|SW_{2}^{2}(f_{\phi}\sharp F,G)-SW_{2}^{2}(f_{\phi^{\prime}}\sharp F,G)\big|\leq\gamma(\|\phi-\phi^{\prime}\|_{2}),$ (30)
for a modulus of continuity $\gamma:\mathbb{R}_{+}\to\mathbb{R}_{+}$ such that $\lim_{t\to 0}\gamma(t)=0$.
Proof.
From Lemma 1, we first bound the difference $|SW_{2}(f_{\phi}\sharp F,G)-SW_{2}(f_{\phi^{\prime}}\sharp F,G)\big|$ for $\phi,\phi^{\prime}\in\Phi$. Using the triangle inequality of $SW_{2}$, we have:
$|SW_{2}(f_{\phi}\sharp F,G)-SW_{2}(f_{\phi^{\prime}}\sharp F,G)|$
$\leq|SW_{2}(f_{\phi}\sharp F,f_{\phi^{\prime}}\sharp F)+SW_{2}(f_{\phi^{\prime}}\sharp F,G)-SW_{2}(f_{\phi^{\prime}}\sharp F,G)|$
$=SW_{2}(f_{\phi}\sharp F,f_{\phi^{\prime}}\sharp F)\ \leq\ W_{2}(f_{\phi}\sharp F,f_{\phi^{\prime}}\sharp F).$ (31)
By definition of Wasserstein distance,
$W_{2}(f_{\phi}\sharp F,f_{\phi^{\prime}}\sharp F)$ $=\inf_{\pi\in\Pi(F,F)}\left(\mathbb{E}_{(X,Y)\sim\pi}\|f_{\phi}(X)-f_{\phi^{\prime}}(Y)\|_{2}^{2}\right)^{1/2}$
$\leq\left(\mathbb{E}_{(X,Y)\sim(Id,Id)\sharp F}\|f_{\phi}(X)-f_{\phi^{\prime}}(Y)\|_{2}^{2}\right)^{1/2}$
$=\left(\mathbb{E}_{X\sim F}\|f_{\phi}(X)-f_{\phi^{\prime}}(X)\|_{2}^{2}\right)^{1/2}$
$\leq\sqrt{\omega(\|\phi-\phi^{\prime}\|_{2}^{2})}$ (32)
due to Assumption 5 Combining with (A.1), we obtain
$\big|SW_{2}(f_{\phi}\sharp F,G)-SW_{2}(f_{\phi^{\prime}}\sharp F,G)\big|\ \leq\ \sqrt{\omega(\|\phi-\phi^{\prime}\|_{2}^{2})}.$ (33)
We now extend (33) to $SW_{2}^{2}$. By the factorization
$\big|a^{2}-b^{2}\big|=|a-b|\,(a+b),$
for $a>0,b>0$,
$\big|SW_{2}^{2}(f_{\phi}\sharp F,G)-SW_{2}^{2}(f_{\phi^{\prime}}\sharp F,G)\big|$
$\leq\big(SW_{2}(f_{\phi}\sharp F,G)-SW_{2}(f_{\phi^{\prime}}\sharp F,G)\big)\big(SW_{2}(f_{\phi}\sharp F,G)+SW_{2}(f_{\phi^{\prime}}\sharp F,G)\big)$
$\quad\leq\sqrt{\omega(\|\phi-\phi^{\prime}\|_{2}^{2})}\,\big(SW_{2}(f_{\phi}\sharp F,G)+SW_{2}(f_{\phi^{\prime}}\sharp F,G)\big).$ (34)
Using $SW_{2}\leq W_{2}$ and the standard bound
$W_{2}(f_{\phi}\sharp F,G)=\sqrt{\inf_{\pi\in\Pi(F,G}\mathbb{E}_{(X,Y)\sim\pi}[\|f_{\phi}(X)-Y\|_{2}^{2}]}$ (35)
$\leq\sqrt{\inf_{\pi\in\Pi(F,G)}\mathbb{E}_{(X,Y)\sim\pi}[\|f_{\phi}(X)\|_{2}^{2}+\|Y\|_{2}^{2}]}$ (36)
$\leq\sqrt{\mathbb{E}_{X\sim F}\|f_{\phi}(X)\|_{2}^{2}}+\sqrt{\mathbb{E}_{Y\sim G}\|Y\|_{2}^{2}},$ (37)
due to the inequality $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$. Together with Assumption 3 and Assumption 5 yield
$SW_{2}(f_{\phi}\sharp F,G)+SW_{2}(f_{\phi^{\prime}}\sharp F,G)\leq\sqrt{C}+\sqrt{M_{2}}$ (38)
Hence,
$\big|SW_{2}^{2}(f_{\phi}\sharp F,G)-SW_{2}^{2}(f_{\phi^{\prime}}\sharp F,G)\big|$ $\leq 2(\sqrt{C}+\sqrt{M_{2}})\sqrt{\omega(\|\phi-\phi^{\prime}\|_{2}^{2})}$ (39)
$:=\gamma(\|\phi-\phi^{\prime}\|_{2}),$ (40)
where $\gamma(t)=2(\sqrt{C}+\sqrt{M_{2}})\sqrt{\omega(t^{2})}$. We can verify that $\lim_{t\to 0}\gamma(t)=0$. We complete the proof of the Lemma. ∎
Lemma 3 (Uniform Law of Large Numbers).
Under assumptions 2,  3, and 5,
$\sup_{\phi\in\Phi}|R_{N}(\phi)-R(\phi)|\overset{a.s}{\to}0\quad\text{as }N\to\infty.$ (41)
Proof.
Since $\Phi$ is compact by Assumption 2, there exists a $\delta$-net $\{\phi_{1},\ldots,\phi_{K}\}\subset\Phi$ such that for any $\phi\in\Phi$, there exists $\phi^{(k)}\in\{\phi_{1},\ldots,\phi_{K}\}$ with
$\|\phi-\phi^{(k)}\|\leq\delta.$ (42)
Using Lemma 2, we have for each $i=1,\ldots,N$:
$|SW_{2}^{2}(f_{\phi}\sharp F_{i},G_{i})-SW_{2}^{2}(f_{\phi^{(k)}}\sharp F_{i},G_{i})|\leq\gamma(\delta),$ (43)
where $\lim_{\delta\to 0}\gamma(\delta)=0$. Averaging over $i$ and taking expectation, this implies
$\frac{1}{N}\sum_{i=1}^{N}|SW_{2}^{2}(f_{\phi}\sharp F_{i},G_{i})-SW_{2}^{2}(f_{\phi^{(k)}}\sharp F_{i},G_{i})|\leq\gamma(\delta),$ (44)
$\mathbb{E}[|SW_{2}^{2}(f_{\phi}\sharp F_{i},G_{i})-SW_{2}^{2}(f_{\phi^{(k)}}\sharp F_{i},G_{i})|]\leq\gamma(\delta)$ (45)
By Jensen’s inequality, we have:
$\left|\frac{1}{N}\sum_{i=1}^{N}SW_{2}^{2}(f_{\phi}\sharp F_{i},G_{i})-\frac{1}{N}\sum_{i=1}^{N}SW_{2}^{2}(f_{\phi^{(k)}}\sharp F_{i},G_{i})\right|\leq\gamma(\delta),$ (46)
$\left|\mathbb{E}[SW_{2}^{2}(f_{\phi}\sharp F_{i},G_{i})]-\mathbb{E}[SW_{2}^{2}(f_{\phi^{(k)}}\sharp F_{i},G_{i})\right|]\leq\gamma(\delta),$ (47)
which is equivalent to
$|R_{N}(\phi)-R_{N}(\phi^{(k)})|\leq\gamma(\delta),\quad|R(\phi)-R(\phi^{(k)})|\leq\gamma(\delta).$ (48)
Using triangle inequality, we have:
$\sup_{\phi\in\Phi}|R_{N}(\phi)-R(\phi)|$
$\leq\sup_{\phi\in\Phi}\left(|R_{N}(\phi)-R_{N}(\phi^{(k)}))|+|R_{N}(\phi^{(k)})-R(\phi)|\right)$
$\leq\sup_{\phi\in\Phi}\left(|R_{N}(\phi)-R_{N}(\phi^{(k)}))|+|R_{N}(\phi^{(k)})-R(\phi^{(k)})|+|R(\phi^{(k)})-R(\phi)|\right)$
$\leq\max_{k=1,\ldots,K}|R_{N}(\phi^{(k)})-R(\phi^{(k)})|+2\gamma(\delta).$ (49)
Let $Z_{N}\;:=\;\sup_{\phi\in\Phi}\,|R_{N}(\phi)-R(\phi)|$ and $A_{N}^{\varepsilon}\;:=\;\{\,Z_{N}>\varepsilon\,\},\ \ \varepsilon>0.$ Almost sure convergence $Z_{N}\overset{a.s}{\to}0$ is equivalent to
$\forall\,\varepsilon>0:\quad\mathbb{P}\!\left(\limsup_{N\to\infty}A_{N}^{\varepsilon}\right)=0,$ (50)
where $\limsup_{N\to\infty}A_{N}^{\varepsilon}:=\bigcap_{m=1}^{\infty}\bigcup_{N\geq m}A_{N}^{\varepsilon}$. For a fixed $\varepsilon>0$, we choose $\delta>0$ so small that $2\gamma(\delta)<\varepsilon/2$. Then (49) implies the event inclusion
$A_{N}^{\varepsilon}\;\subseteq\;\bigcup_{k=1}^{K}\left\{\,|R_{N}(\phi^{(k)})-R(\phi^{(k)})|>\frac{\varepsilon}{2}\right\}=:\bigcup_{k=1}^{K}B_{N,k}^{\varepsilon/2}.$ (51)
Therefore,
$\limsup_{N\to\infty}A_{N}^{\varepsilon}\;\subseteq\;\bigcup_{k=1}^{K}\left(\limsup_{N\to\infty}B_{N,k}^{\varepsilon/2}\right).$ (52)
By the strong law of large numbers (using Assumption 3) applied to each fixed $\phi^{(k)}$,
$|R_{N}(\phi^{(k)})-R(\phi^{(k)})|\overset{a.s}{\to}0,$ (53)
hence, for every $k$ and every $\eta>0$,
$\mathbb{P}\!\left(\limsup_{N\to\infty}\{\,|R_{N}(\phi^{(k)})-R(\phi^{(k)})|>\eta\,\}\right)=0.$ (54)
Taking $\eta=\varepsilon/2$ and using the finite union bound, we get
$\mathbb{P}\!\left(\limsup_{N\to\infty}A_{N}^{\varepsilon}\right)\;\leq\;\sum_{k=1}^{K}\mathbb{P}\!\left(\limsup_{N\to\infty}B_{N,k}^{\varepsilon/2}\right)=0.$ (55)
Since this holds for every $\varepsilon>0$, it follows that
$\forall\,\varepsilon>0:\quad\mathbb{P}\!\left(\limsup_{N\to\infty}A_{N}^{\varepsilon}\right)=0,$ (56)
which is equivalent to $Z_{N}\to 0$ almost surely, i.e.,
$\sup_{\phi\in\Phi}|R_{N}(\phi)-R(\phi)|\overset{a.s}{\to}0.$ (57)
Back to the main proof of posterior consistency, for all $\epsilon>0$, we define the “bad” set
$S_{\epsilon}(\phi_{0}):=\{\phi\in\Phi:\|\phi-\phi_{0}\|_{2}\geq\epsilon\},$ (58)
where $\phi_{0}$ is defined in Assumption 1. We want to show
$\pi_{N}(S_{\epsilon}(\phi_{0}))\overset{a.s}{\to}0,$ (59)
for all $\epsilon>0$. For $\varepsilon>0$, let
$A_{N}:=\big\{\pi_{N}(S_{\epsilon}(\phi_{0}))>\varepsilon\big\}.$ (60)
By the definition of almost sure convergence in terms of the limit superior of events, it suffices to show
$\mathbb{P}\Big(\limsup_{N\to\infty}A_{N}\Big)=0,$ (61)
where $\limsup_{N\to\infty}A_{N}:=\bigcap_{m=1}^{\infty}\bigcup_{N\geq m}A_{N}.$ By the uniform law of large numbers (Lemma 3), for any $\eta>0$ define the event
$E_{N}:=\Big\{\sup_{\phi\in\Phi}|R_{N}(\phi)-R(\phi)|\leq\eta\Big\}.$ (62)
Then $E_{N}$ occurs eventually almost surely. On $E_{N}$:
  • For $\phi\in S_{\epsilon}(\phi_{0})$,
    $R_{N}(\phi)\geq R(\phi)-\eta\geq R(\phi_{0})+\Delta-\eta,$ (63)
    where $\Delta=\Delta(\epsilon)\;=\;\inf_{\{\phi\in\Phi:\|\phi-\phi_{0}\|_{2}\geq\epsilon\}}\big(R(\phi)-R(\phi_{0})\big)\;>\;0$ (Assumption 1).
  • For $\phi\in B_{\delta}(\phi_{0}):=\{\phi:\|\phi-\phi_{0}\|_{2}<\delta\}$, choose $\delta$ small so that
    $\sup_{\phi\in B_{\delta}(\phi_{0})}(R(\phi)-R(\phi_{0}))\leq\eta,$ (64)
    then
    $R_{N}(\phi)\leq R(\phi)+\eta\leq R(\phi_{0})+2\eta.$ (65)
On $E_{N}$, the posterior mass ratio
$\frac{\pi_{N}(S_{\epsilon}(\phi_{0}))}{\pi_{N}(B_{\delta}(\phi_{0}))}$ $=\frac{\int_{S_{\epsilon}(\phi_{0})}\exp(-wNR_{N}(\phi))\,d\pi(\phi)}{\int_{B_{\delta}(\phi_{0})}\exp(-wNR_{N}(\phi))\,d\pi(\phi)}$
$\leq\frac{\pi(\Phi)\,\exp(-wN(R(\phi_{0})+\Delta-\eta))}{\pi(B_{\delta}(\phi_{0}))\,\exp(-wN(R(\phi_{0})+2\eta))}$
$=\frac{1}{\pi(B_{\delta}(\phi_{0}))}\exp(-wN(\Delta-3\eta)),$ (66)
which implies
$\pi_{N}(S_{\epsilon}(\phi_{0}))\leq C\exp(-wN(\Delta-3\eta)),$ (67)
where $C=\frac{1}{\pi(B_{\delta}(\phi_{0}))}>0$ due to Assumption 4. Choose $\eta<\Delta/4$ so that $\Delta-3\eta>\Delta/4>0$., then the right-hand side decays exponentially in $N$, implying that for almost every sample path $\omega$ there exists $N_{0}(\omega)$ such that for all $N\geq N_{0}(\omega)$,
$\pi_{N}(S_{\epsilon}(\phi_{0}))<\varepsilon.$ (68)
By the definition of $\limsup$ of events,
$\limsup_{N\to\infty}A_{N}=\bigcap_{m=1}^{\infty}\bigcup_{N\geq m}\{\pi_{N}(S_{\epsilon}(\phi_{0}))>\varepsilon\}=\emptyset\quad\text{a.s.}$ (69)
Hence,
$\mathbb{P}\Big(\limsup_{N\to\infty}A_{N}\Big)=0.$ (70)
Since the choice of $\varepsilon>0$ is arbitrary, we conclude
$\pi_{N}(S_{\epsilon}(\phi_{0}))\xrightarrow{\text{a.s.}}0.$ (71)
Thus the posterior concentrates around $\phi_{0}$ for any $\epsilon>0$, proving Bayesian posterior consistency.
Proposition 1 (Linear function).
Let $f_{\phi}(x)=Ax+b$ ($\phi=(A,b)$), the following holds:
$\mathbb{E}_{X\sim F}\|f_{\phi}(X)-f_{\phi^{\prime}}(X)\|_{2}^{2}\leq(M_{1}+1)\|\phi-\phi^{\prime}\|_{2}^{2},$ (72)
for all $(F,G)\in\mathrm{supp}(P)$, $\phi,\phi^{\prime}\in\Phi$, and $M_{1}>0$ in Assumption 3. In addition, for all $(F,G)\in\mathrm{supp}(P)$ and $\phi\in\Phi$,
$\mathbb{E}_{X\sim F}[\|f_{\phi}(X)\|_{2}^{2}]\leq C,$ (73)
for a constant $C<\infty$.
Proof.
Using the Cauchy–Schwarz inequality,
$\mathbb{E}_{X\sim F}[\|f_{\phi}(X)-f_{\phi^{\prime}}(X)\|_{2}^{2}]$ $=\mathbb{E}_{X\sim F}[\|AX+b-A^{\prime}X+b^{\prime}\|_{2}^{2}]$
$=\mathbb{E}_{X\sim F}[\|(A-A^{\prime})X+(b-b^{\prime})\|_{2}^{2}]$
$\leq\mathbb{E}_{X\sim F}[(\|A-A^{\prime}\|_{F}^{2}+\|b-b^{\prime}\|_{2}^{2})(\|X\|_{2}^{2}+1)]$
$\leq(\|A-A^{\prime}\|_{F}^{2}+\|b-b^{\prime}\|_{2}^{2})(\mathbb{E}_{X\sim F}[\|X\|_{2}^{2}+1])$
$=(\|A-A^{\prime}\|_{F}^{2}+\|b-b^{\prime}\|_{2}^{2})(M_{1}+1)$
$=(M_{1}+1)\|\phi-\phi^{\prime}\|_{2}^{2}.$ (74)
due to Assumption 3.
We further have:
$\mathbb{E}_{X\sim F}[\|f_{\phi}(x)\|_{2}^{2}]$ $=\mathbb{E}_{X\sim F}[\|AX+b\|_{2}^{2}]$
$\leq\mathbb{E}_{X\sim F}[(\|A\|_{F}^{2}+\|b\|_{2}^{2})(\|X\|_{2}^{2}+1)]$ (75)
$\leq(\|A\|_{F}^{2}+\|b\|_{2}^{2})(M_{1}+1),$ (76)
due to Assumption 3. Since the parameter space $\Phi$ is compact (Assumption 2), $(\|A\|_{F}^{2}+\|b\|_{2}^{2})$ is bounded, hence,
$\mathbb{E}_{X\sim F}[\|f_{\phi}(x)\|_{2}^{2}]\leq C,$ (77)
for a constant $C=(\|A\|_{F}^{2}+\|b\|_{2}^{2})(M_{1}+1)>0$, which completes the proof. ∎
Appendix B MCMC